What will be the probability of 4 sons to a couple?
Since the probability of having a son is about 1/2, the probability of the first 4 children being boys is about (1/2)4.
You mean the PROBABILITY?
It's 13/52*13/51 = 13/204
Is it possible to have an event that causes more than one hazardous process to occur?
It is possible to have an event that causes more than one hazardous process to occur. An example is the landslide in Myanmar. Because of the heavy rains, a flood and landslide can both occur. Another example can be seen in the game Sim City. If there is an earthquake in the town, the earthquake can cause fires by the power lines.
How do you apply maths in your daily life?
Maths is your whole day ... :)) u see time calculate time it includes additions subtractions .... u write days expenditure in your notebook for example then it involves more complex maths ....
It wont be enough to say how we use maths in daily life ..
Updated answer
If you use the tripartite system you will see simple maths that we use in our every day lives. We see positive and negative things as we go through our day, mostly we draw conclusions and move on. These conclusions are the same as central tendency. The mean the mode and the median are used to find an average or draw conclusions for or against general problems.
In a tripartite way we live in the 1)now and draw conclusions from our 2) past experience and us these conclusions in the 3) future.
An interesting second paragraph - a concept new to me, but I assume taken from statistics!
It's really an open-ended question, as "everyday life" can range from arithmetic as simple as tour suggested adding your shopping bill, via the usual mensuration & trig used in the building & engineering trades, to the genuinely complex maths used in engineering-design, science etc.
Differences between probability samplingand non-probability sampling?
the difference is just that non-probability sampling does not involve random selection, but probability sampling does.
How probability useful in whether forecasting?
Most people are not statistically trained so the probability of whether or not you are forecasting is so close to 0 that knowing its value is of little help.
What is the main advantage of a multistage rocket?
that the total weight of the rocket is greatly reduced as the rocket rises
What is the probability of getting exactly six tails if you flip a coin eight times?
(8!/(6!2!) / 2**8 = 28/256 = 7/64 = 0.109375
What is the probability of rolling 5 on one die three times in a row?
In the long run, it is a certainty.
In the first three throws it is 1/216.
There are 9+6 = 15 checkers in the bag. 6 of them are red. 6 out of 15 are red. Drawing a red checker has a probability of P = 6/15 = 2/5 = 0.4 = 40%
Since you replace the checker, the probability Q that red is drawn again remains 0.4.
The probability of both events occurring (red drawn twice) equals the product of probabilities, PQ = (0.4)*(0.4) = 0.16.
If a die is rolled once what is the probability of getting more than 3?
There is a 1 in 2 chance, so 1/2.
How many ways can you roll a five with 2 dice or greater?
6 ways: 1 and 4, 4 and 1, 2 and 3, 3 and 2, 5 and 0, 0 and 5
A five can be rolled with either dice being a 2,3,4,1 and the other dice being the correlating number that adds to five ie. 4-1, 1-4. 3-2. 2-3.
There is no 0 on a die.
What does it means if the standard deviation is large?
that you have a large variance in the population and/or your sample size is too small
What is the exact range of Hounsfield unit?
; Hounsfield unit is a unit used in medical imaging (MSCT ) to describe the amount of x-ray attenuation of each "voxel" in the 3D image. The voxels are normally represented as 12-bit binary numbers, and therefore have 212 = 4096 possible values. These values are arranged on a scale from -1024 HU to +3071 HU, calibrated so that -1024 HU is the attentuation produced by air and 0 HU is the attenuation produced by water.
What is the probability of male child when a baby is born?
According to the American CIA, worldwide the number of boys born to the number of girls is 1.07:1, which implies a probability of about 0.517. This varies somewhat by country, for instance. At one extreme, according to the list of countries by sex ratio presented in wikipedia is Liechtenstein at 0.558, at the other Haiti at 0.503.